OpenStudy (anonymous):

y=1/5 x+6 what is the ordered pair?

OpenStudy (anonymous):

determine if (5,31) is a solution

OpenStudy (anonymous):

sustitute x in the equation and check if y = the solutiom

OpenStudy (anonymous):

whether (5.31) is a solution that is easy you just need to replace x by 5 and then see if the resulting y is 31

OpenStudy (anonymous):

I get yes it is cos 5x5=25+6=31, right?

OpenStudy (anonymous):

what is the equation????

OpenStudy (anonymous):

huh@josefelk, shr mnt cos=because

OpenStudy (anonymous):

you can rewrite it

OpenStudy (anonymous):

am i right though

OpenStudy (anonymous):

yes u are

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

y=1/5 x+6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I wrote it kind of crazy, sorry

OpenStudy (anonymous):

it does not matter